prove that force equals the rate of change of momentum
1
Expert's answer
2019-06-07T11:44:29-0400

From second Newtons law:


"m\vec{a}=m\frac{d\vec{v}}{dt}=\vec{F}=\frac{d(m\vec{v})}{dt}=\frac{d\vec{p}}{dt}"